OpenStudy (anonymous):

Chris went to the store to purchase 5.5 gallons of milk. When he got to the store, he noticed that the milk was no longer sold in gallons but in liters. Which of the following shows the number of liters, with the correct number of significant digits, Chris needs to buy in order to have 5.5 gallons of milk? (1 liter = 0.26418 gallon)

OpenStudy (jhannybean):

\[\ 5.5 \ gal \cdot \frac{1 \ L}{0.26418 \ gal}\]\[\ 5.5 \ \cancel{gal} \cdot \frac{1 \ L}{0.26418 \ \cancel{gal}}\]

OpenStudy (jhannybean):

\[\ \frac{5.5 \ L}{0.26418}\] will give you the number of L